OpenStudy (anonymous):

Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3 Just having some troubles need help @hero

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

My function is f(x)=29 (6%)x

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

I think you got it right

OpenStudy (anonymous):

I just need to figure how to graph them wit my function

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I have f(x + 2)=29(6%)^x for the first one if you can show me how to graph the first one I should be fine with the rest.